OpenStudy (anonymous):

Determine which polynomial is a perfect square trinomial. 25x^2 − 40x − 16 9a^2 − 20a − 25 25b^2 − 15b + 9 16x^2 − 56x + 49

OpenStudy (cwrw238):

note it cant be the first or second choice because there are no real square roots of -16 or -25

OpenStudy (anonymous):

well ow is it one of the last two? i cant find one that works

OpenStudy (cwrw238):

oh sorry - my mistake - you have to multiply them then add take the the third one: 25b^2 − 15b + 9 you have sqrt25 = 5 and sqrt 9 = -3 ( i took negative sqrt becuse of the -15x) so that 5 * -3 = - 15 and -15 + -15 = -30 - so cant be this one

OpenStudy (anonymous):

okay, so it would be the last one?

OpenStudy (cwrw238):

4th option we got 16x^2 − 56x + 49 sqrt 16 = 4 and sqrt 49 = -7 so 4 * -7 = -28 and -28 + -28 = -56 we got the one

OpenStudy (anonymous):

okay thank you

OpenStudy (cwrw238):

you can check it out by the distributive law; (4x - 7)(4x -7 ) = 16x^2 -28x - 28x + 49 = 16x^2 - 56x + 49

OpenStudy (cwrw238):

yw